OpenStudy (anonymous):

let f(x) = x^2-3 and g(x)=3x+1. Find (fog)(x)

OpenStudy (anonymous):

(fog)(x) = f(g(x)

OpenStudy (anonymous):

ie = f (3x+1)

OpenStudy (anonymous):

= (3x+1)^2 - 3

OpenStudy (anonymous):

Solve this..)

OpenStudy (anonymous):

\[(a+b)^2 = a^2 + 2ab + b^2\]

OpenStudy (anonymous):

does this then equal 9x^2 -2?

OpenStudy (anonymous):

u missed to use 2ab

OpenStudy (anonymous):

9x^2+6x-2

OpenStudy (anonymous):

Is this the correct answer then 9x^2+6x-2?

OpenStudy (anonymous):

f(x) = x^2-3 and g(x)=3x+1 f(g(x)) = (3x+1)^2 - 3 = 9x^2 + 6x + 1 - 3 = 9x^2 + 6x - 2 yes.